# Task and Motion Planning in Hierarchical 3D Scene Graphs

### v1

@@ Abstract
Recent work in the construction of 3D scene graphs has enabled mobile robots to build large-scale hybrid metric-semantic hierarchical representations of the world. These detailed models contain information that is useful for planning, however how to derive a planning domain from a 3D scene graph that enables efficient computation of executable plans is an open question. In this work, we present a novel approach for defining and solving Task and Motion Planning problems in large-scale environments using hierarchical 3D scene graphs. We identify a method for building sparse problem domains which enable scaling to large scenes, and propose a technique for incrementally adding objects to that domain during planning time to avoid wasting computation on irrelevant elements of the scene graph. We test our approach in two hand crafted domains as well as two scene graphs built from perception, including one constructed from the KITTI dataset. A video supplement is available at https://youtu.be/63xuCCaN0I4.
I INTRODUCTION
@@
We aim to enable an autonomous agent to solve large-scale Task and Motion Planning (TAMP) problems in real-world environments. In order to solve a TAMP problem, an abstract planning domain is needed, which efficiently and accurately represents the robot’s environment and the available actions. Over the past several years, significant progress has been made in the area of generating hybrid metric-semantic representations of the world using 3D scene graphs. One such example is the Hydra scene graph [12], which jointly represents low-level information such as scene geometry and robot trajectories, and higher level information such as semantically-annotated object instances, places, rooms, and buildings. This environmental abstraction lends itself well to large-scale TAMP problems, as it is capable of storing both higher level abstractions like objects and connectivity of regions needed for task planning, as well as the low-level metric information required to check kinematic feasibility of different actions [12]. However, how best to construct a planning domain from the 3D scene graph representation remains an open question.
The trend in scene graph development has been toward encoding an increasing amount of information in the hierarchical representations. Such detailed information may be crucial to the success of a planner, as the more faithfully any planning domain (regardless of the planning system or formalism) represents the real world, the more likely a found solution will actually be executable by the robot. Moreover, increased representational capacity of the planning domain leads to a wider range of robot goals that can be specified, thus increasing flexibility and generality.
@@
However, as a planning problem instance grows in the number of objects, so too does the complexity of finding a plan. TAMP is PSPACE-Hard [22], so problems can become computationally intractable very quickly as the sizes of the state and action spaces grow [7]. To avoid creating intractable planning problems when converting a 3D scene graph into a planning problem, it is critical to determine which elements of the environment (i.e., which elements of the scene graph) are relevant. Consider, for example, a robot responding to a Chemical, Biological, Radiological, Nuclear, and Explosive (CBRNE) scenario, receiving instructions to inspect and potentially neutralize dangerous objects scattered in a large area, represented as a pre-built scene graph. The robot is able to pass near an object only after it has neutralised and cleared it, and may be instructed to avoid particular regions entirely, to solve different tasks. Depending on the geometry of the scene and the specified goal, only a subset of these potentially dangerous obstacles and regions may ultimately be relevant to finding a plan. But, for a robot building a scene graph representation from perception, it is not at all obvious which elements of the scene graph should be added to a planning domain to ensure a valid plan can be found and executed. If we were to add every potentially relevant element to our planning domain, we would be guaranteed to eventually find a plan which takes our robot to its goal, but at the expense of substantial computational complexity. On the other hand, if we are too aggressive in pruning away seemingly irrelevant elements, we may restrict ourselves to sub-optimal solutions, or fail to find a plan at all.
@@
Previous approaches to the problem of inferring a task-relevant planning domain have relied on representations of connectivity in the scene graph to prune superfluous elements [1]. However, these efforts have been limited to very specific task planning problems, as the pruning approaches employed often remove information necessary for checking the geometric feasibility of plans, or implicitly limit the types of goals that can be specified. Alternative approaches for reducing the planning domain size involve attempting to learn the relevance of planning objects, then incrementally adding objects to the domain according to the learned relevance score until the problem is solvable [18]. Unfortunately, this approach requires training on numerous similar planning problems, and is difficult to generalize to tasks at large scales.
In this work, we propose a novel approach to both enable and accelerate TAMP in large environments (Fig. 1). The first contribution of this work is the definition of a set of properties that a TAMP domain must possess to ensure that any plans returned by solving a problem actually meet the user’s specifications. We then define a method for translating a Hydra scene graph [12] into a planning domain that meets these criteria. Our second contribution is the formulation of a sufficient condition for removing symbols from a planning problem while maintaining feasibility. This condition shows that, for example, many of the places in a complete scene graph can be ignored when formulating certain kinds of planning problems, greatly reducing planning times. This improvement relies on the definition of a hierarchical planner, which ignores certain elements of the scene at the highest level to accelerate discrete planning, while planning over the entire scene at lower levels to ensure constraints on the plan are not violated. The third contribution is a method to further accelerate planning by incrementally identifying objects in the scene as relevant during search according to how they affect the feasibility of certain abstract plans. In contrast to the places component of the domain, finding a small set of objects ahead of time to include in the planning instance while guaranteeing that the problem remains feasible is quite difficult. Instead, each time we fail to solve a sub-problem in an abstract plan, we use the scene graph to identify which objects were responsible. We then augment our domain with those objects, and continue planning. We show the effectiveness of our approach across two hand-crafted domains, in addition to two scene graphs built from real perception.
II Task and Motion Planning in 3D Scene Graphs
@@
Figure 1: An illustration of how we derive our planning representation from a 3D scene graph generated from the KITTI dataset. In (A), we present three levels of the scene graph hierarchy, with the low-level Mesh layer on the bottom, the Places layer above that, and the Objects layer on the top. In (B), we see an isometric view of this graph, giving an insight to the scale, while also plotting the higher level Regions layer. In (C) we show a very simplified version of this scene, where the agent is given a task specified in logic. The agent is tasked with either visiting Place 6 while avoiding Place 1, or visiting Place 5. Given the geometry of the scene, the first clause in this disjunction is un-satisfiable. We also see that Place 5 is partially obstructed by a suspicious object, so the agent must consider either avoiding it (green trajectory), or inspect and neutralize the object (blue trajectory) before continuing to its goal. To solve problems like this one at large scales, we take advantage of the structure of the scene graph to accelerate planning.
TAMP jointly considers elements of high-level task planning [8, 13] and low-level motion planning [15] in an attempt to solve hybrid discrete/continuous, multi-modal planning problems [7].
II-A Task and Motion Planning Preliminaries
@@
A common formalism for encoding planning problems is Planning Domain Definition Language (PDDL). In a PDDL problem, a state $\mathcal{I}$ is a set of facts: instances of boolean functions called predicates $p(\bar{x})\in\mathcal{P}$, which are parameterized by tuples of objects $\bar{x}=[x_{1},...,x_{k}]$, $x\in\mathcal{O}$. Transitions between states are defined by actions $a(\bar{x})\in\mathcal{A}$ (also parameterized by objects) which are expressed as two sets of predicates: preconditions $\text{Pre}(a_{i})$ and effects $\text{Eff}(a_{i})$. An action’s preconditions determine if an instance can be applied from a particular state, and its effects define the set of facts that are added ($\text{Eff}^{+}(a_{i}))$ or removed $(\text{Eff}^{-}(a_{i}))$ from the state. A planning domain is composed of lifted sets of predicates $\mathcal{P}$ and actions $\mathcal{A}$, and a problem instance $P=(\mathcal{P},\mathcal{A},\mathcal{O},\mathcal{I}_{0},\mathcal{G})$ combines a domain with an initial state $\mathcal{I}_{0}$ and a set of goal states $\mathcal{G}$, parameterized by objects $\mathcal{O}$. We will refer to objects in a PDDL problem as symbols when they may be confused with physical objects from a scene graph.
One approach for representing a TAMP problem —which we use in this work— is an extension of PDDL called PDDLStream [6]. A PDDLStream problem instance $(\mathcal{P},\mathcal{A},\mathcal{S},\mathcal{O},\mathcal{I}_{0},\mathcal{G})$ represents the discrete search portion of a TAMP problem in PDDL, with the same representation of predicates, actions, objects, initial state, and set of goal states. In a TAMP problem however, certain actions may have symbol parameterisations that are either cumbersome or intractable to add to a problem instance. For example, consider a move action, which defines robot motion from one configuration to another, and which may have a precondition that requires the existence of a collision-free trajectory between those configurations. Adding all such trajectory symbols between all possible points in the configuration space to the planning domain is impossible, so instead, we define sub-problems called streams $s\in\mathcal{S}$ which can be solved to return symbols that (if a solution exists) can be used to certify the preconditions of the associated action. In the case of a trajectory, the relevant sub-problem/stream might be a motion planner, which, if solved, returns a symbol representing a collision-free trajectory between two configurations that the robot can follow. We refer the reader to [6] for a detailed description of PDDLstream.
@@
Solutions to PDDLStream problems take the form of a sequence of parameterized action instances $\pi=[a_{1}(\bar{x}_{1}),a_{2}(\bar{x}_{2}),...,a_{n}(\bar{x}_{n})]$ that define a plan, where assigned parameters $\bar{x}$ satisfy each action’s constraints [7]. From the parameters of an action, we can derive a motion sequence, which specifies how a robot executes a particular action. For example, from an action plan composed of a sequence of move actions, the corresponding motion sequence would be composed of the trajectories that were solved for by the motion-planning sub-problem and which are parameters of those move actions. Executing that sequence would involve multiple calls to a trajectory controller. We say that two motion sequences are equivalent if they result in the agent acting identically. For any action sequence, there is a corresponding sequence of states $\mathcal{I}_{\pi}=[\mathcal{I}_{0},\mathcal{I}_{1},\mathcal{I}_{2},...,\mathcal{I}_{n}]$, leading from the initial state to a goal state that can be constructed from each action’s effects. For an action plan $\pi$, its corresponding state plan $\mathcal{I}_{\pi}$ is valid if $\mathcal{I}_{i}\in\text{Pre}(a_{i+1})$ for $i=0,...,N-1$, and $\mathcal{I}_{N}\in\mathcal{G}$. Tasks specified in PDDL can be solved by a range of solvers [9, 11], and any state plan found by such a solver is valid by construction. A feasible planning problem is one for which there is a valid solution.
II-B Building 3D Scene Graphs from Perception
In most TAMP settings, the planning domain and problem instance are assumed to be given ahead of time. However, a robot operating in the real world must be able to build its own world model from perception, especially the symbols and predicates that describe the world states. Recently, 3D scene graphs such as those implemented by Hydra [12] or S-Graphs [3] have emerged as promising hierarchical representations that can be built in real-time onboard robots, to provide abstractions that are useful for encoding planning domains.
@@ II-B Building 3D Scene Graphs from Perception
Our definition of a 3D scene graph, directly based on Hydra  [12], consists of several layers of increasing abstraction (see Fig. 1). Each layer consists of a collection of nodes representing the location and other attributes of an element, with edges connecting nodes within the same layer representing relative spatial constraints and edges between different layers representing an inclusion relationship. The lowest layer of the hierarchy is a semantically-annotated mesh of the scene geometry. The next layer contains objects identified by a semantic image segmentation. The places layer represents navigable regions of the environment based on semantic and geometric properties of the mesh. Places are clustered into groups based on geometric and semantic information, and these groups become nodes in the higher-level regions layer (e.g., rooms in an indoor environment).
Previous work on 3D scene graphs has mainly focused on indoor uses. These representations rely on the Generalized Voronoi Diagram (GVD) [16] to generate places, an abstraction of 3D spatial connectivity, which are not well suited for ground robot navigation. We use an alternate formulation of 2D places in our navigable scene graph. Each of these places represents a 2D polygon with consistent terrain classification, and is connected to adjacent places in the scene graph. As the resolution of the places is much coarser than the mesh resolution, planning over sequences of places can be much faster than planning over the mesh itself, while still retaining important geometric information. The learned metric-semantic method for extracting regions from places presented in Strader et al. [21] is more appropriate for these 2D places compared to the purely geometric method room-based region extraction originally presented in Hughes et al. [12]. Hydra and its region-based clustering extension [21] can construct this map representation in realtime from RGBD sensor data while accounting for odometry drift, enabling large scale, consistent, and information-rich maps.
II-C Inferring the Planning Domain from Scene Graphs
@@
We introduce a method for deriving a planning problem instance from a Hydra scene graph to demonstrate the salient aspects of solving planning problems based on large-scale environments. In general, we generate six classes of predicates that may be relevant for scene graph related planning. 1) Type information derived from the nodes of the graph, where each node is a unary predicate: (Configuration ?c), (Place ?p), and (Object ?o), etc. 2) Agent/object predicates that define the state of the robot and objects: (AtConfig ?c), (AtPlace ?p), (AtRoom ?r), etc. 3) Connection predicates defined by edges in the same level of the graph: (Connected ?n1 ?n2). 4) Hierarchical predicates indicating edges connecting nodes of different levels of the graph: (PoseInPlace ?c ?p), (PlaceInRoom ?p ?r), etc. 5) Preconditions of actions that are certified by solving a stream’s associated sub-problem. For the example move action, the predicate certified by the motion-planner would contain the two configurations, and the trajectory between them: (Trajectory ?c1 ?t ?c2). 6) Additional, problem specific predicates defined by the user that can be used to specify goal states and problem constraints such as which places to visit or which objects to collect.

### v2

# MIT, Cambridge, MA 02142

@@ Abstract
Recent work in the construction of 3D scene graphs has enabled mobile robots to build large-scale metric-semantic hierarchical representations of the world. These detailed models contain information that is useful for planning, however an open question is how to derive a planning domain from a 3D scene graph that enables efficient computation of executable plans. In this work, we present a novel approach for defining and solving Task and Motion Planning problems in large-scale environments using hierarchical 3D scene graphs. We describe a method for building sparse problem instances which enables scaling planning to large scenes, and we propose a technique for incrementally adding objects to that domain during planning time that minimizes computation on irrelevant elements of the scene graph. We evaluate our approach in two real scene graphs built from perception, including one constructed from the KITTI dataset. Furthermore, we demonstrate our approach in the real world, building our representation, planning in it, and executing those plans on a real robotic mobile manipulator. A video supplement is available at https://youtu.be/v8fkwLjBn58.
1 Introduction
@@
We aim to enable an autonomous agent to solve large-scale Task and Motion Planning (TAMP) problems in real-world environments. In order to do so efficiently, an abstract planning domain is needed, which accurately represents the robot’s environment as well as its available actions. Recently, significant progress has been made in the area of generating hierarchical metric-semantic representations of the world using 3D scene graphs [15, 4]. These environmental abstractions lend themselves well to large-scale planning problems, as they are capable of storing both higher-level abstractions such as objects and connectivity of regions which are needed for task planning, as well as the low-level metric information required to check kinematic feasibility of different actions.
@@
However, as a planning problem instance grows in the number of objects, so too does the computational burden of finding a plan. TAMP is PSPACE-Hard [27], so problems can become computationally intractable very quickly as the sizes of the state and action spaces grow [10]. To create tractable planning problems when converting a 3D scene graph into a planning domain, it is critical to leverage the scene graph’s structure and identify which elements of the environment are relevant. Consider, for example, a robot responding to a Chemical, Biological, Radiological, Nuclear, and Explosive (CBRNE) scenario, receiving instructions to inspect and neutralize dangerous objects scattered in a large area, represented as a scene graph. The robot can pass near an object only after it has neutralized and cleared it, and the robot may be instructed to avoid particular regions entirely. Depending on the geometry of the scene and the specified goal, only a subset of these dangerous obstacles and regions may ultimately be relevant to finding a plan. But, for a robot building a scene graph representation from perception, it is not at all obvious which elements of the scene graph should be added to a planning domain to ensure a valid plan can be found and executed.
@@
Previous approaches to the problem of inferring a task-relevant planning domain have relied on representations of connectivity in the scene graph to prune superfluous elements [1]. However, these efforts have been limited to specific kinds of task planning problems, as the pruning approaches employed often remove information necessary for checking the geometric feasibility of plans, or they implicitly limit the types of goals that can be specified. Alternative approaches for reducing the planning problem size involve attempting to learn the relevance of planning objects, then incrementally adding objects to the domain according to the learned relevance score until the problem is solvable [24]. Unfortunately, this approach requires training on numerous similar planning problems, and is difficult to generalize to tasks at large scales in the real world.
@@
Figure 1: An illustration of how we derive and encode tasks in our planning representation from a 3D scene graph. (A) An isometric view of a Hydra scene graph generated from the KITTI dataset, giving an insight to the scale of the environment. (B) A simplified version of this scene, where the agent is tasked with either visiting Place 6 while avoiding Place 1, or visiting Place 5. We see that Place 5 is partially obstructed by a suspicious object, so the agent must consider either avoiding it (green trajectory), or inspecting and neutralizing the object (blue trajectory) to reach its goal. (C) A mobile robot (which we used to build our scene graphs) executing a plan in the real world, inspecting an object in the top frame, and moving an obstruction out of its path on the bottom.
In this work, we propose a novel approach to both enable and accelerate TAMP in large environments (Fig. 1). Our first contribution is a three-level hierarchical planner for planning in large domains derived from 3D scene graphs. Next, we present the formulation of a sufficient condition for removing symbols from a planning problem while maintaining feasibility, which can greatly reduce computation when planning. This condition shows that many of the places in a 3D scene graph can be ignored when formulating planning problems that factorize according to our three-level hierarchy. We introduce a technique for reasoning about whether the sparsified domain matches the original intent of the planning domain, which reveals extra constraints that must be imposed on the motion planner. Finally, we develop a method to further accelerate planning by incrementally identifying objects in the scene as relevant during search according to how the geometry of the scene affects the feasibility of certain high-level plans. We show the effectiveness of our approach across two hand-crafted domains, two scene graphs built from real perception with planning in simulation, and finally a real-world mobile manipulation task on a Spot robot.
2 Task and Motion Planning in 3D Scene Graphs
In this section, we introduce how we encode TAMP problems, review the 3D scene graph structure that we leverage for grounding planning problems, and finally propose a CBRNE-inspired planning domain as an example of formulating a planning problem based on a 3D scene graph.
2.1 Task and Motion Planning Preliminaries
@@
The TAMP problem jointly considers elements of high-level task planning [11, 16] and low-level motion planning [20] in an attempt to solve hybrid discrete/continuous, multi-modal planning problems [10]. A common formalism for encoding task planning problems is the Planning Domain Definition Language (PDDL). In a PDDL problem, a state $\mathcal{I}$ is a set of facts, where each fact is an instance of a boolean function called a predicate $p(\bar{x})\in\mathcal{P}$, which is parameterized by a tuple of symbols $\bar{x}=[x_{1},\ldots,x_{k}]$ from a given set of symbols $\mathcal{O}=\{x_{i}\}$. Each symbol $x_{i}$ is a discretization of a state variable. Transitions between states are defined by actions $a(\bar{x})\in\mathcal{A}$ (also parameterized by symbols) which are expressed as two sets of predicates: preconditions $\text{Pre}(a_{i})$ and effects $\text{Eff}(a_{i})$. An action’s preconditions determine if an instance can be applied from a particular state $\mathcal{I}$, and its effects define the set of facts that are added ($\text{Eff}^{+}(a_{i}))$ or removed $(\text{Eff}^{-}(a_{i}))$ from the state $\mathcal{I}$. A planning domain is composed of lifted sets of predicates $\mathcal{P}$ and actions $\mathcal{A}$, and a problem instance $P=(\mathcal{P},\mathcal{A},\mathcal{O},\mathcal{I}_{0},\mathcal{G})$ combines a domain with an initial state $\mathcal{I}_{0}$ and a set of goal states $\mathcal{G}$, parameterized by symbols $\mathcal{O}$.
@@
Solutions to PDDL problems take the form of a sequence of parameterized action instances $\pi=[a_{1}(\bar{x}_{1}),a_{2}(\bar{x}_{2}),...,a_{n}(\bar{x}_{n})]$ where the state after taking each action satisfies the precondition of the following action [10]. For any action sequence, there is a corresponding sequence of states $\mathcal{I}_{\pi}=[\mathcal{I}_{0},\mathcal{I}_{1},\mathcal{I}_{2},...,\mathcal{I}_{n}]$, leading from the initial state to a goal state that can be constructed from each action’s effects. We will use the fact that only a subset of state symbols are needed for each action to enable a factorization of the planning problem in Section 3.1. For an action plan $\pi$, its corresponding state plan $\mathcal{I}_{\pi}$ is valid if $\mathcal{I}_{i}\in\text{Pre}(a_{i+1})$ for $i=0,...,N-1$, and $\mathcal{I}_{N}\in\mathcal{G}$. A range of solvers [13, 14] can solve tasks specified in PDDL, and any state plan found by such a solver is valid by construction. A feasible planning problem is one for which there is a valid solution.
The continuous nature of TAMP problems, coupled with the scale of environments we consider here, make discretizing and encoding a planning problem directly in pure PDDL infeasible. We therefore use an extension of PDDL called PDDLStream [9] to represent and solve TAMP problems. A PDDLStream problem instance $(\mathcal{P},\mathcal{A},\mathcal{S},\mathcal{O},\mathcal{I}_{0},\mathcal{G})$ represents the discrete search portion of a TAMP problem in PDDL, as a set of predicates, actions, symbols, initial state, and set of goal states, but also introduces the notion of streams $s\in\mathcal{S}$, which can query external samplers/solvers (e.g. a motion planner) during search to produce new symbols and facts within the problem instance. Streams make the problem encoding more efficient, as they obviate the need to evaluate predicates for all possible continuous values of a symbol. PDDLStream solves [^1] Specifically, the PDDLStream adaptive solution algorithm. problems by first finding an optimistic solution that satisfies the domain’s symbolic constraints – a task skeleton – and then attempting to solve for feasible continuous parameters. We refer the reader to [9] for a detailed description of PDDLstream.
In a TAMP problem, each symbol $x_{i}$ can represent a continuous value (e.g., a pose), and the grounded parameters of an action depend on these values. From these parameters, we can derive a motion sequence, which specifies how a robot executes an action. For example, from an action plan composed of a sequence of move actions, the corresponding motion sequence would be composed of the trajectories that were solved for by the motion planner and describes the continuous values of the parameters $\bar{x}_{i}$ for each move action. Executing that sequence involves multiple calls to a trajectory controller, where two motion sequences are equivalent if they result in the agent acting identically.
2.2 Building 3D Scene Graphs from Perception
@@
A robot operating in the real world ideally is able to build its own PDDL domain model from perception. We assume the robot is equipped with a prior model of its own state and a motion controller, as well as the ability to use its sensors to build a dense geometric model of its environment and the objects in it. To derive a discrete, symbolic model of the geometry, we take advantage of recent work in 3D scene graph mapping [15] that infers a discretization of the geometry and the objects in the geometric model. While our approach is compatible with a range of scene graph implementations, our definition of a 3D scene graph, directly based on Hydra  [15, 26], consists of several layers of increasing abstraction (see Fig. 1). Each layer consists of a collection of nodes representing location and other attributes, with edges connecting nodes within the same layer representing relative spatial constraints and edges between different layers representing an inclusion relationship. The lowest layer of the hierarchy is a semantically-annotated mesh of the scene geometry. The next layer contains objects and their locations identified by a semantic image segmentation. The places layer represents navigable regions of the environment based on semantic and geometric properties of the mesh. Places are clustered into groups based on geometric and semantic information, and these groups become nodes in the higher-level regions layer (e.g., rooms in an indoor environment). Hydra can construct this map representation in realtime from RGBD sensor data while accounting for odometry drift, enabling large scale, consistent, and information-rich maps.
Previous work on 3D scene graphs has mainly focused on indoor uses. These representations rely on the Generalized Voronoi Diagram (GVD) [22] to generate places, an abstraction of 3D spatial connectivity, which are not well suited for ground robot navigation. We use an alternate formulation of 2D places in our navigable scene graph, where each place represents a 2D polygon with consistent terrain classification, representing an area the robot may traverse (Fig. 1B).
2.3 Inferring the Planning Domain from Scene Graphs
@@
We introduce a framework for deriving a TAMP problem instance from a scene graph to demonstrate the salient aspects of solving planning problems based on large-scale environments. In general, the problem contains a symbol $x\in\mathcal{O}$ for each node in the scene graph, as well as a symbol corresponding to the robot. We define six classes of predicates, derivable from a Hydra scene graph, that may be relevant for planning: 1) Type information derived from the nodes of the graph, where each node corresponds to a unary predicate: (Configuration ?c), (Place ?p), and (Object ?o), etc. 2) Agent or object predicates that define the state of the robot and objects: (AtConfig ?c), (AtPlace ?p), (AtRoom ?r), etc. 3) Connection predicates defined by edges in the same level of the graph: (Connected ?n1 ?n2), 4) Inclusion predicates indicating edges connecting nodes of different levels of the graph: (PoseInPlace ?c ?p), (PlaceInRoom ?p ?r), etc. 5) Preconditions of actions that are certified by solving a stream’s associated sub-problem. For example, a move action may require that a trajectory has been found between two configurations: (Trajectory ?c1 ?t ?c2). 6) Additional, problem-specific predicates defined by the user to specify goal states and problem constraints such as which places to visit or which objects to collect.